# Detecting the hidden population of low-mass haloes in strong lenses

### v1

@@ 4.1 Effect of lens galaxy angular structure
We find that the addition of multipoles to the lens galaxy model greatly reduces the detectability of the population of low-mass haloes. For the most extreme model, with multipoles <3<3 per cent, the effect is completely destroyed. In O’R24 we trained models $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, also used here, with circular multipoles. The deflection angle due to the circular multipoles is trivial to calculate, but this formulation adds perturbations in a circular basis to a mass distribution in an elliptical basis, and so for lenses which are far from circular, the mass distribution can become unrealistic.
@@
Since these models were trained a computationally friendly calculation of the deflection angle in the elliptical basis has been found (Paugnat & Gilman, 2025). In this basis, the angular perturbations are less localised and more closely follow the macro structure of the main lens galaxy. As Paugnat & Gilman (2025) show, elliptical multipoles produce smaller flux ratio anamolies in simulated lensed quasars than circular multipoles (Cohen et al., 2024). It should then be expected that the elliptical multipole perturbations are less degenerate with an individual low-mass halo, although this has not yet been systematically tested.
@@
The results in Section˜3.5 and the strong dependence on the angular structure allowed in the lens galaxy, however formulated, again underline the need for a comprehensive study of the angular mass-density structure in real lens galaxies. Although the use of elliptical multipoles is well motivated by the isophotes of ellitpical galaxies, it is yet unclear whether the isophotes are an indicator of the mass distribution to the accuracy required for dark matter substructure studies. For now, the results in Fig.˜7 and Section˜3.5 can be thought of as presenting the most pessimistic case for low-mass halo population detection, with the reality likely lying somewhere between the $\mathcal{M}_{1}$ (Fig.˜4) and $\mathcal{M}_{2}$ cases.
@@ 4.2 Prospects for dark matter inference
We have shown that strong lens observations are sensitive to perturbations from populations of individually undetectable subhaloes. This population exists in a mass range which carries a large amount of information about the dark matter model, information which is presently discarded by the sensitivity map. In CDM, this population of objects with masses $M_{\mathrm{max}}>10^{6}M_{\odot}$ exists in abundance in every strong lens galaxy, so that studies which find an excess of detections in many lenses would obtain strong evidence in favour of CDM. In WDM models, this population is suppressed for masses $M_{\mathrm{max}}<M_{\mathrm{hm}}$. Finding no excess of detections in many lenses would be evidence in favour of a value of $M_{\mathrm{hm}}$ above the lower mass limit of the population. In Fig.˜5 we show this to be up two orders of magnitude below the individual sensitivity limit. Although the constraints in Fig.˜6 are conservative, as they do not make use of mass and concentration information, taking account of the population effect would significantly improve the state of the art constraints on dark matter from strong lensing (Enzi et al., 2021).
@@ 5 Conclusions
    We find a $40$ per cent excess in the number of detections in CDM, i.e., our subhalo detector finds subhaloes with $1.4$ times the frequency predicted by the sensitivity map. This excess dissapears in warmer dark matter models, where the population of subhaloes is suppressed at the low mass end.
@@
    Finally, we repeat our analyses for subhalo detection models which include circular multipole perturbations to the lens galaxy. We find that the multipoles present a strong degeneracy with the signal of this low-mass population, much more so than for individually detectable objects. However, this result may change if elliptical multipoles were used, which may be less degenerate with the subahlo signal in general.
Future dark matter studies with strong lensing have the capability to dramatically increase their constraining power by properly accounting for this effect, especially for studies using large sky surveys like Euclid, Roman, and SKA. However, as with individual detections of subhaloes, an understanding of the mass structure of the lens galaxy is critical for avoiding systematic biases and incorrect inferences.

### v2

@@ 4.1 Systematic uncertainty due to lens galaxy properties
We find that the addition of large amplitude, circular multipoles to the lens galaxy model greatly reduces the detectability of the population of low-mass haloes. For the most extreme model, with multipoles <3<3 per cent, the effect is completely destroyed. In O’R24 we trained models $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, also used here, with circular multipoles. The deflection angle due to the circular multipoles is trivial to calculate, but this formulation adds perturbations in a circular basis to a mass distribution in an elliptical basis, and so for lenses which are far from circular, the mass distribution can become unrealistic.
@@
Since these models were trained a computationally friendly calculation of the deflection angle in the elliptical basis has been found (Paugnat & Gilman, 2025). In this basis, the angular perturbations are less localised and more closely follow the macro structure of the main lens galaxy. As Paugnat & Gilman (2025) show, elliptical multipoles produce smaller flux ratio anomalies in simulated lensed quasars than circular multipoles (Cohen et al., 2024). It should then be expected that the elliptical multipole perturbations are less degenerate with an individual low-mass halo, although this has not yet been systematically tested.
@@
The results in section˜3.5 and the strong dependence on the angular structure allowed in the lens galaxy, however formulated, again underline the need for a comprehensive study of the angular mass-density structure in real lens galaxies. Although the use of elliptical multipoles is well motivated by the isophotes of elliptical galaxies, it is yet unclear whether the isophotes are an indicator of the mass distribution to the accuracy required for dark matter substructure studies. Therefore, the results in fig.˜7 and section˜3.5 can be thought of as presenting the worst-case scenario for low-mass halo population detection. This experiment would ideally be repeated using elliptical multipoles and a prior informed by measured multipoles in the total mass profiles of real strong lenses, or at least in simulated galaxies. Such measurements are currently too few in number to build an accurate prior. In such a test, it is likely that the result would lie somewhere between the $\mathcal{M}_{1}$ (fig.˜4) and $\mathcal{M}_{2}$ cases, presenting more optimistic prospects for constraining the low-mass halo population.
As well as the lens galaxy’s angular structure, its emission can also interfere with the detection of dark substructures if not properly dealt with (e.g. Nightingale et al., 2024). In this work, we assumed that the lens light is perfectly modelled and subtracted before the data are analysed for substructures. This is realistic for radio observations, where the lens galaxy typically has no emission (e.g. Powell et al., 2025) but for optical and NIR wavelengths, the lens galaxy light presents an issue. When detecting a population of substructures, whose effect is more diffuse than that of a single perturber, the degeneracy with the lens galaxy light is likely to be stronger.
@@ 4.2 Prospects for dark matter inference
We have shown that strong lens observations are sensitive to perturbations from populations of individually undetectable subhaloes. This population exists in a mass range which carries a large amount of information about the dark matter model, information which is presently discarded by the sensitivity map. In CDM, this population of objects with masses $M_{\mathrm{max}}>10^{6}M_{\odot}$ exists in abundance in every strong lens galaxy, so that studies which find an excess of detections in many lenses would obtain strong evidence in favour of CDM. In WDM models, this population is suppressed for masses $M_{\mathrm{max}}<M_{\mathrm{hm}}$. Finding no excess of detections in many lenses would be evidence in favour of a value of $M_{\mathrm{hm}}$ above the lower mass limit of the population. In fig.˜5 we show this to be up two orders of magnitude below the individual sensitivity limit. Although the constraints in fig.˜6 are conservative, as they do not make use of mass and concentration information, taking account of the population effect would significantly improve the state of the art constraints on dark matter from strong lensing (Enzi et al., 2021). Scatter in the subhalo mass-concentration relation also means that CDM models produce some subhaloes with higher than average concentration. This further increases the observable difference between CDM and WDM, where these outliers are absent (Amorisco et al., 2022). In this work we required the properties of the simulated subhalo populations and the subhaloes used in the sensitivity map to correspond exactly for the same mass. As such, we could not account for the effect of this scatter on the dark matter constraints.
@@ 5 Conclusions
    We find a $40$ per cent excess in the number of detections in CDM, i.e., our subhalo detector finds subhaloes with $1.4$ times the frequency predicted by the sensitivity map. This excess dissipaters in warmer dark matter models, where the population of subhaloes is suppressed at the low mass end.
@@
    Finally, we repeat our analyses for subhalo detection models which include varying amplitudes of circular multipole perturbations to the lens galaxy. We find that the multipoles present a strong degeneracy with the signal of this low-mass population, much more so than for individually detectable objects. However, this result would change if more physically motivated elliptical multipoles were used, which may be less degenerate with the subhalo signal in general.
Future dark matter studies with strong lensing have the capability to increase their constraining power by properly accounting for this effect, especially for studies using large sky surveys like Euclid, Roman, and SKA. However, as with individual detections of subhaloes, an understanding of the mass structure of the lens galaxy is critical for avoiding systematic biases and incorrect inferences.